### v1

# Study of symmetries in finite temperature $N_{f}=2$ QCD with Möbius Domain Wall Fermions

Abstract
We report on the ongoing study of symmetry of $N_{f}=2$ QCD around the critical temperature. Our simulations of $N_{f}=2$ QCD employ the Möbius domain-wall fermion action with residual mass $\sim 1~{}\rm{MeV}$ or less, maintaining a good chiral symmetry. Using the screening masses from the two point spatial correlators we compare the mass difference between channels connected through various symmetry transformations. Our analysis focuses on restoration of the $SU(2)_{L}\times SU(2)_{R}$ as well as anomalously broken axial $U(1)_{A}$. We also present additional study of a potential $SU(2)_{CS}$ symmetry which may emerge at sufficiently high temperatures.
1 Introduction
Study of $SU(2)_{L}\times SU(2)_{R}$ and $U(1)_{A}$ (broken spontaneously by quantum anomaly) in $N_{f}=2$ is an important supplement to the study of Quantum Chromodynamics (QCD) with nondegenerate flavor such as the studies done by [1, 2, 3, 4, 5, 6]. The core physics of $SU(N_{f})_{L}\times SU(N_{f})$ primarily relies on $N_{f}=2$ as all nondegenerate flavors have large masses and will explicitly break the $SU(N_{f})$ symmetry down to the two lowest mass quarks. Therefore, direct study of the chiral phase transition in 2-Flavor QCD is quite straightforward with a core physical signature which is replicated in QCD with larger flavor symmetry. Chiral symmetry breaking is of interest to ongoing work concerning the phases of QCD, as it is broken spontaneously at low temperatures and may be related to higher temperature symmetries associated with deconfinement through the chiral condensate. It has also been argued that the chiral condensate is not simply an order parameter for chiral symmetry transitions, but may in fact have a relationship to the susceptibility to the axial anomaly $U(1)_{A}$ which is broken anomalously below the scale of QCD. While the anomaly has been considered a low energy effect, there is some evidence[7, 8, 9, 10, 11, 12] that the chiral condensate is connected to reduced susceptibility to the quantum anomaly at high temperatures. The relationship of these two symmetries, and their transitions around the critical point are important to the continued study of high temperature QCD and deconfined matter which were present at early times in the universe.
In addition to the well known symmetries of QCD, we are interested in studying additional emergent symmetry structures which are not seen in the original QCD Lagrangian. In recent literature [13, 14, 15, 16, 17] emergent higher symmetries of QCD e.g. $SU(N_{f})_{CS}$, $SU(2N_{f})$ have been increasingly discussed, which we have addressed in previous numerical simulations [18, 19, 3, 20, 21, 22]. In particular JLQCD has simulated $N_{f}=2$ QCD and focused on the temperatures above the phase transition $T\sim 165-330~{}\rm{MeV}$ using Möbius Domain Wall Fermions. In this work, we extend this study to include simulations just below the critical temperature, as well as, including the psudocritical temperature $T_{c}=165(3)~{}\rm{MeV}$ predicted from chiral susceptibiltiy in [23, 24, 25, 26]. In addition to this we make use of the spatial two point correlation functions to extract screening masses of mesons in $I=1$ channel. For various choices of $\gamma_{k}$ we can use channels of the spatial correlations functions to study $SU(2)_{L}\times SU(2)_{R}$, $U(1)_{A}$ and $SU(2)_{CS}$ by way of direct mass difference between the channels related by various symmetry transformations.
2 Möbius Domain Wall Fermions
The four dimensional fermion action is described by an approximation of the sign function in the four dimensional overlap operator, we will use the form of Nueberger’s overlap operator:
$D_{OV}=\frac{1+m}{2}-\frac{1-m}{2}\gamma_{5}\mbox{sgn}(H).$ (1)
Our choice of Möbius domain wall fermions is related to our choice in the form of the kernel function $H$ which is
$H=\frac{\gamma_{5}D_{W}}{2+\gamma_{5}D_{W}}$
For a sufficiently large but finite fifth dimensional extent $L_{s}$ we can approximate the function $\mbox{sgn}(H)$ as $\tanh(H)$ like under the assumption that the eigenvalue spectrum of the kernel $H$ is of $\mathcal{O}(1)$.
$\epsilon(H)=\dfrac{(H+1)^{L_{s}}-(H-1)^{L_{s}}}{(H+1)^{L_{s}}+(H-1)^{L_{s}}}=\tanh(L_{s}\tanh^{-1}(H)).$ (2)
For our study we take $L_{s}=16$ which reduces the residual mass or violation of the Ginsparg-Wilson relation to $<1~{}\rm{MeV}$. This minimizes the symmetry breaking effects for screening masses for $\mathcal{O}(100~{}\rm{MeV})$ and above, assuring that all of our meson correlators are chiral at the temperatures in our range of study. The key benefit for this work is that we have a theoretically clean approach to analysis of correlation functions because symmetry transformations applied to each channel can be studied from the data directly [27, 28, 29].
3 Mesonic Correlators
3.1 Spatial Mesonic Correlators
Let us consider the quark isospin triplet bilinear operators
$\mathcal{O}_{\Gamma}(x)=\overline{q}(x)(\Gamma\otimes\frac{\vec{\tau}}{2})q(x),$ (3)
where $\tau^{a}$ is an element of the generators of $SU(2)$, $\Gamma$ denotes the combination of Dirac $\gamma$ listed in Table. 1. Using the correlation functions of the $\mathcal{O}_{\Gamma}(x)$ operators we examine symmeries at high temperature using the screening mass spectrum. Specifically we make use of the two point correlator along the $z$-axis,
$C_{\Gamma}(n_{z})=\sum_{n_{t},n_{x},n_{y}}\braket{\mathcal{O}_{\Gamma}(n_{x},n_{y},n_{z},n_{t})\mathcal{O}^{\dagger}_{\Gamma}(0,0,0,0)}$ (4)
were $n_{t}$ is the Euclidean time direction and we specify the correlator channel with $\Gamma$. At long distances the correlation function is expected to exponentially decay following $C_{\Gamma}(z)\sim e^{-M_{\Gamma}z}$, where $M_{\Gamma}$ is the screening mass of the lowest energy state in the $\Gamma$ channel.
As the quarks are subject to anti-periodic boundary conditions in the temporal direction, the momentum recieves a large contribution from the Matsubara frequencies. If the largest contribution to the two-point correlation function is from a pair of static non-interacting quarks, then the screening mass $M_{\Gamma}$ is well approxmiated to $2\pi T$, twice the lowest Matsubara frequency. Therefore, a comparison between these two quanitites $M_{\Gamma}$ and $2\pi T$ provides an effective estimation of the strength of interaction between quarks in QCD.
$\Gamma$ Reference Name Abbr. Symmetry Correspondences
$\mathbb{I}$ Scalar S }$\bigg{\}}U(1)_{A}$
$\gamma_{5}$ Psuedo Scalar PS
$\gamma_{k}$ Vector $\mathbf{V}_{k}$$\bigg{\}}SU(2)_{L}\times SU(2)_{R}$$\bigg{\}}SU(2)_{CS}?$
$\gamma_{k}\gamma_{5}$ Axial Vector $\mathbf{A}_{k}$
$\gamma_{k}\gamma_{3}$ Tensor $\mathbf{T}_{k}$$\bigg{\}}U(1)_{A}$
$\gamma_{k}\gamma_{3}\gamma_{5}$ Axial Tensor $\mathbf{X}_{k}$
Table 1: Here we express the set of observables associated with the quark correlation functions, and the symmetry group transformations which connect the various operators. In this table we have chosen to project the tensor channels onto states in the $z$-direction.
3.2 Emergence of $SU(2)_{CS}$
For $T\rightarrow\infty$ the free quark propagator appears to exhibit an additional set of symmetries. Let us consider a quark propagator which has momenta $(p_{x},p_{y})$ perpendicular to the $z$-axis:
$\braket{\overline{q}(z)q(0)}(p_{1},p_{2})=\sum_{p_{0}}\int_{-\infty}^{\infty}\frac{dp_{3}}{2\pi}\dfrac{m-(i\gamma_{0}p_{0}+i\gamma_{i}p_{i})}{p_{0}^{2}+p_{i}^{2}+m^{2}}e^{ip_{3}z},$ (5)
where we have decomposed the Euclidean time direction into the Matsubara frequencies $p_{0}=2\pi T(n+\frac{1}{2})$ and labelled $x,y$ and $z$ momenta with numerical indicies. We can integrate spatially by taking the $z$-directional momentum corresponding to the “energy spectrum” $E=\sqrt{p_{0}^{2}+m^{2}+p_{1}^{2}+p_{2}^{2}}$,
$\braket{\overline{q}(z)q(0)}(p_{1},p_{2})=\sum_{p_{0}}\dfrac{m+\gamma_{3}E-(i\gamma_{0}p_{0}+i\gamma_{1}p_{1}+i\gamma_{2}p_{2})}{2E}e^{-Ez}.$ (6)
In the $T\rightarrow\infty$ limit we can take $T\gg m^{2}+p_{1}^{2}+p_{2}^{2}$ and expand the correlator in terms of $1/T$:
$\braket{\overline{q}(z)q(0)}=\gamma_{3}\frac{1+i\mbox{sgn}(p_{0})\gamma_{0}\gamma_{3}}{2}e^{-\pi Tz}+O(1/T).$ (7)
Here we have approximated the summation over $p_{0}$ to be the lowest Matsubara mode only. This form of the correlator is invariant under an additional set of transformations:
$\begin{split}q(x)&\rightarrow\exp\bigg{(}{\frac{i}{2}\Sigma_{a}\theta_{a}}\bigg{)}q(x)\\ \overline{q}(x)&\rightarrow\overline{q}(x)\gamma_{0}\exp\bigg{(}-\frac{i}{2}\Sigma_{a}\theta_{a}\bigg{)}\gamma_{0},\end{split}$ (8)
where the set of transformations is defined by
$\Sigma_{k}=\begin{bmatrix}\gamma_{k}\\ i\gamma_{k}\gamma_{5}\\ \gamma_{5}\end{bmatrix}.$
The set of the generators of the so-called $SU(2)_{CS}$ chiral-spin or $SU(2)_{CS}$ symmetry, explored in [17, 16, 18, 30, 31]. However, for the case of QCD with quarks interacting with the gluonic medium the symmetry does not appear as directly as the free quark example above, and appears to be approximate. This is due to contributions from the chromoelectic term of the quark-gluon interaction(when fixed to the rest frame of the medium), as this term is symmetric under the same transformations $\Sigma$. However, the symmetry breaks down at higher temperatures due the increase in chromomagnetic effects and thus the emergent symmetry of the chromoelectic and quarks field breaks in the high temperature limit[30, 1, 2, 5].
4 Numerical Results
Our simulations of $N_{f}=2$ flavor QCD employ the tree-level improved Symanzik gauge action, as well as Möbius domain wall fermion action with stout smearing. For all ensembles $\beta=4.30$ with a lattice cutoff set to $a^{-1}=2.463~{}\rm{GeV}$. The residual mass of the domain-wall quark is reduced to $<1~{}\rm{MeV}$. For the highest temperature ensembles the aspect ratio is 4:1 and we do expect finite size effects on the screening masses. For temperatures $T\sim 147~{}\rm{MeV}$ and $T\sim 165~{}\rm{MeV}$ lattices with spatial extent $L=36$ and $L=32$ respectively may not be enough to control the long range correlation effects for the spatial correlation functions and so we also consider additional lattices with spatial extents of $L=48$ and $L=40$ respectively. For all new ensemble data we considered four values of $am=0.0010,0.0025,0.00375,0.0050$ for up and down quark masses with the lightest value $am=0.0010$ corresponding to a mass of $2.6~{}\rm{MeV}$ below the physical point for the lightest quarks. Masses above $13~{}\rm{MeV}$ corresponding to $am=0.0050$, were simulated but not analyzed as part of the set of temperature dependent screening masses; for completeness these will be listed in forthcoming publication of our results.
The value for the screening mass is extracted from fitting lattice data to the standard form
$C(z)=A\cosh(m[z-L/2]).$ (9)
Previous work on $N_{f}=2$ QCD have also considered an $C(z)\sim\exp(-mz)/z$ fitting form, which is predicted from the free two quark propagator; however, for the low temperatures in this study, the free quark fitting ansatz cannot be applied.
Figure 1 is an example of effective mass for the $PS$ 1(a) and $T$ 1(b) channels, which are taken from our lightest mass $m=0.0010$ ensemble at temperature $T=147~{}\rm{MeV}$. For fittings which have significant noise, both the effective mass fits as well as correlator fits were used to determine a reliable fitting range.
Figure 1: Above are the effective mass plots for the $PS$ and $Tt$ channels with $m=0.0010$ at $T=147~{}\rm{MeV}$.
Figure 2: For the lightest mass quark ensembles the $N_{f}=2$ screening mass data plotted with respect to temperature, shown as points. The shaded bands are $N_{f}=2+1$ data from HotQCD in [32]. For $N_{f}=2$ there is great overlap with the $T=0$ meson spectrum at $T=0.9T_{c}$. In the high temperature regime both the $N_{f}=2$ and $N_{f}=2+1$ converge to twice the lowest Matsubara frequency plotted in grey. The $X$ and $T$ tensor channels rapidly converge to $2\pi T$ after the chiral crossover point.
Figure 3: This plot is a version of fig.2 with both the $N_{f}=2$ and $N_{f}=2+1$ normalized by their respective $T_{c}$ to create a common transition point. The screening masses are also normalized by the lowest order Matsubara frequency $2\pi T$. The agreement between $N_{f}=2$ and $N_{f}=2+1$ data is excellent in this normalized plot.
Figure 4: Above is the screening mass difference plot for $SU(2)_{L}\times SU(2)_{R}$ with respect to temperature for the range of masses in this study. Restoration is quite clean and exact for the lowest mass $am=0.0010$.
Figure 5: Above is the screening mass difference plot for $U(1)_{A}$ with respect to temperature, susceptibility to the axial anomaly appears to be minimal at $T=185~{}\rm{MeV}$.
Figure 6: Above is the screening mass difference plot for $SU(2)_{CS}$ for various temperatures and masses, while it looks close to being restored at $T=330~{}\rm{MeV}$ it is quite far from “restoration”.
Figure 7: Multiple quark masses are plotted for the highest temperature lattice $T=330~{}\rm{MeV}$, for the all quarks $SU(2)_{CS}$ remains broken.
5 Summary of Preliminary Results
In figure 2 the fitted mass versus temperature is plotted for the lightest mass quarks. Due to the difference in the psuedocritical transition temperature for $N_{f}=2$ and $N_{f}=2+1$ QCD figure 3 is the plot of temperature normalized by $T_{c}$ and fitted masses normalized by $2\pi T$. This allows a direct comparison between the 2019 results from HotQCD in [32] and our 2-flavor study.
Figure 2 shows, for the lowest temperature in the study, a large splitting between correlator channels related by symmetry transformation consistent with broken chiral symmetry and susceptibility to the axial anomaly. At $T=147~{}\rm{MeV}$ we see very good overlap between screening mass in various channels and the low temperature meson bound state masses; in particular, the psuedoscalar overlaps with the $T=0$ $\pi^{\pm}$ mass at $0.9T_{c}$ and $T_{c}$ indicating that the chiral condensate strongly correlates with the broken chiral symmetry.
The most significant difference between our work and the results by HotQCD, which are plotted as shaded bands in figs 2 and 3, is the behavior of the $S$ channel at low temperatures. Because our simulations are done with Möbius Domain Wall fermions we do not observe the same $2\pi^{\pm}$ lattice artifact, and instead, we observe a large increase in the screening mass in the region of broken chiral symmetry.
$SU(2)_{L}\times SU(2)_{R}$ symmetry restoration, seen through the mass difference in the $V$ and $A$ partner channels, shows a very clear restoration signal for the all masses but is then most evident in the $am=0.0010$. From figure 4 at $T=147~{}\rm{MeV}$ chiral symmetry is broken, and it appears to be restored at the transition point $T_{c}\sim 165~{}\rm{MeV}$ and remains zero for all temperatures. $U(1)_{A}$ symmetry is also effectively “restored” for $T\sim 165~{}\rm{MeV}$ from the plot of $X_{t}-T_{t}$ fig. 5. While it is still unclear if the transition is exactly at the psuedocritical point at $T=188~{}\rm{MeV}$ corresponding to $T\sim 1.1T_{c}$ we see restoration of $U(1)_{A}$ which is below the $1.2-1.3T_{c}$ threshold for restoration reported by [32, 25, 33, 34, 35, 12] and is observed in the $PS-S$ screening mass difference as well.
In contrast to both of these symmetries we also plotted a temperature spectrum for the $SU(2)_{CS}$ predicted in [15, 16, 17], unlike the clear signals of symmetry restoration seen in figs 4 and 5, the plot of the symmetry pair $X-A$ figure 5 does not converge to zero for the temperatures in our study. While there is a significant reduction in noise consistent with the symmetry restoration for the other symmetry pairs, we do not see a convergence to, or sustained, zero value indicating emergence of $SU(2)_{CS}$. This is seen also in figure 7 which shows $SU(2)_{CS}$ in comparison to other symmetries at $T=330~{}\rm{MeV}$ for which there is not even an approximate threshold as the channels connected by $SU(2)_{CS}$ remain significantly broken, which is true for all temperatures in this study.
Acknowledgments
We thank L. Glozman and Y. Sumino for useful discussions. We would also like to thank O.Philipsen and R.Pisarski for their discussion about our results. For the numerical simulation we use the QCD software packages Grid [36, 37] for configuration generations and Bridge++ [38, 39] for measurements. Numerical simulations were performed on Wisteria/BDEC-01 at JCAHPC under a support of the HPCI System Research Projects (Project ID: hp170061) and Fugaku computer provided by the RIKEN Center for Computational Science under a support of the HPCI System Research Projects (Project ID: hp210231). This work is supported in part by the Japanese Grant-in-Aid for Scientific Research (No. JP26247043, JP18H01216, JP18H04484, JP18H05236,JP22H01219) and by Joint Institute for Computational Fundamental Science (JICFuS).